The interest rate on a $100,000 loan is 7.5% compounded quarterly. What quarterly payments will reduce the balance to $75,000 after five years? (Do not round intermediate calculations and round your final answer to 2 decimal places.)

PMT $
Answer #1

Remaining balance after 5 years or 5*4 = 20 quarters is 75000

Interest rate per quarter is 7.5%/4 = 1.875%

Thus, we have

75000 = A(P/A, 1.875%, 20)

A = 75000(A/P, 1.875%, 20)

= 75000*0.060421

= $4,531.61

Hence, quarterly payments of $4,531.61 will reduce the balance to $75,000 after five years